OpenStudy (anonymous):

solve by factoring 6x^2+22x-40=0

OpenStudy (kamille):

6x^2+22x-40=0 D=22^2-4*(-40)*6=484+960=1444 x1=-22+38/12=1 1/3 x2=-22-38/12=-11/6

OpenStudy (anonymous):

i got -5 and 4

OpenStudy (callisto):

6x^2+22x-40=0 2(3x^2+11x-20)=0 (3x^2+11x-20)=0 (3x-4)(x+5) =0 (3x-4)=0 or (x+5) =0 x = 4/3 or x=-5

OpenStudy (anonymous):

x=-5